OpenStudy (anonymous):

Find the product: -9x(-6x2 + x + 5)

OpenStudy (anonymous):

im assuming to meant 6x^2 \[54x^{3} -9x ^{2}-45x\]

OpenStudy (anonymous):

-9*-6=54 -9*1=-1 -9*5= -45 Yes I did mean -6x^2

OpenStudy (anonymous):

yep..thats correct:) just dont forget to multiply the x as well...

OpenStudy (anonymous):

okay will remember that

OpenStudy (anonymous):

actually: check the second one you wrote: -9*1 = -9...not -1

OpenStudy (anonymous):

yea I noticed that when double checking thanks

OpenStudy (anonymous):

no problem :)